OpenStudy (anonymous):

Using the following equation, find the center and radius: x2 - 4x + y2 + 8y = -4 a. The center is located at (-2, -4), and the radius is 4. b. The center is located at (2, -4), and the radius is 4. c. The center is located at (-2, -4), and the radius is 16. d. The center is located at (2, -4), and the radius is 16.

Directrix (directrix):

You will need to write the equation in this form: (x-h)^2 + (y-k)^2 = r^2 where (h,k) is the center and r is the radius.

OpenStudy (anonymous):

(x-2)^2 + (y+4)^2 = 16 ?

Directrix (directrix):

(x-2)^2 + (y+4)^2 = 4^2 to highlight the radius.

Directrix (directrix):

Compare to this: (x-h)^2 + (y-k)^2 = r^2 where (h,k) is the center and r is the radius.

OpenStudy (anonymous):

.b. The center is located at (2, -4), and the radius is 4.